OpenStudy (anonymous):

Solve the following system of equations involving a circle and a line. x^2+y^2=25 y=x+1

OpenStudy (anonymous):

 just substitute: x^2 + (x+1)^2 = 25 x^2 + x^2 + 2x + 1 = 25 2x^2 + 2x - 24 = 0 x^2 + x - 12 = 0 (x+4)(x-3) = 0 x=-4, x=3 (-4, -3) and (3, 4) are your solutions